Find all the values of the parameter a for which the point of minimum of the function  f(x)=1+a2x-x3 satisfies the inequality ((x2+x+2)/(x2+5x+6))>0

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear harshit

((x2+x+2)/(x2+5x+6))>0

x2+x+2 is always positive

so x2+5x+6 should be >0

 which give

  x<-3  and x>-2

Now

f(x)=1+a2x-x3

  for minimum value differentiate

f'(x)=a2-3x2

  f''(x)=-6x

equate f'(x) =0

  fives x=+-a/√3

case 1

a>0

x=-a/√3 gives minimum value

so    -a/√3 <-3    and   -a/√3 >-2

           a>3√3        and     a<2√3

 so    0<a<2√3   and    a>3√3

 

case 2 

a<0

x=a/√3 gives minimum value

 

so    a/√3 <-3    and   a/√3 >-2

           a<-3√3        and     a>-2√3

 so    a<-3√3   and    -2√3<a<0

 


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin